# No Bells, Just Whistles: Sports Field Registration by Leveraging Geometric Properties

### v2

@@ 3.1.2 Keypoint Disambiguation
Due to the multi-view nature of the SoccerNet dataset [12] and, considering, for instance, one of the soccer field’s semi-circles, ambiguity appears in its respective $\mathcal{K}p_{1}$ and $\mathcal{K}p_{2}$ keypoints candidates, as shown in Fig. 2-bottom. To avoid that, we define two different strategies to handle that disambiguation depending on the total number of keypoints generated in the previous sets: when there are sufficient points in the $\mathcal{K}p\cup\mathcal{K}e$ set to infer a homography, i.e., four points, $\mathcal{K}p_{1}$ is computed first by choosing the candidates combination that minimizes a reprojection error. Then, we include $\mathcal{K}p_{1}$ in the homography estimation and apply the same strategy to $\mathcal{K}p_{2}$. Otherwise, we perform a grid-search involving both $\mathcal{K}p_{1}$ and $\mathcal{K}p_{2}$ candidates when $\mathcal{K}p\cup\mathcal{K}p_{e}\cup\mathcal{K}p_{1}^{*}\cup\mathcal{K}p_{2}^{*}\geq 4$, where * denotes a possible candidate combination. The grid-search iterates over all keypoints candidates in a set-wise manner to avoid unfeasible combinations and keeps the one with minimum reprojection error. It is worth noting that no more keypoint sets beyond $\mathcal{K}p$ are computed if neither of the above stated conditions are met. Moreover, once we define the new sets, two additional constraints are applied in the case homography estimation or ellipse fitting is not accurate enough. Initially, we manually establish a reprojection error threshold to validate points. Subsequently, through iteration over combinations of keypoints, we construct vectors and ensure that the cross-products maintain consistent signs in both world and image coordinates. This final step is essential in cases where two distinct combinations yield valid top- and bottom-view perspectives of the field while exhibiting identical reprojection errors. Utilizing cross-products enables us to differentiate and retain the keypoint combination corresponding to the field’s top-view. The full keypoint generation process is depicted in Fig. 1-top.
3.1.3 Left-Right Disambiguation
In sequences where the camera angle aligns with the longitudinal axis of the court, an ambiguity arises regarding the distinction between the right and left halves of the field. Hence, a critical step to ensure consistency and robustness across keypoints and lines detection processes involves differentiating between the two sides. This is accomplished by implementing a remap to the ground-truth (GT) values, ensuring that the goal area closest to the camera consistently represents the left side. The process to check whether or not the mapping should be applied is defined in a heuristic fashion. We compute angles of horizontal and vertical soccer field lines, respectively, and then set a threshold taking into account angle distribution and visual inspection.
@@ 3.1.3 Left-Right Disambiguation
3.2 Keypoints and Lines Detection
@@
Our approach is built upon Falaleev et al. [15] solution, which makes use of two encoder-decoder convolutional neural networks to estimate the position of the pre-defined keypoints and the soccer field lines depicted in [8] excluding conics, giving the last one an auxiliary role to enhance keypoint completeness. During inference, the former produces heatmaps for each pre-defined keypoint with a single Gaussian peak with $2px$ sigma positioned in the keypoint location, accompanied by an additional target channel. This additional channel reflects the inverse of the maximum value among the other target feature maps, ensuring that the resultant target tensor behaves as a probability distribution function at every spatial point. Meanwhile, the latter network produces heatmaps for each visible soccer field line within the frame, assigning two Gaussian peaks at the line extremities’ locations. Additionally, we introduce an extra channel, known as the boundary channel, to our heatmap following the approach outlined in [35]. This augmentation aims to enhance the efficient capture of global information regarding the soccer field and improve extremities detection, particularly near image borders. We effectively extract the positions of keypoints and line extremities from the generated heatmaps by employing a max pooling operation, drawing inspiration from the methodology proposed in [40]. This process is summarized on Fig. 1-bottom.
3.2.1 Architecture
@@
Following [15], the keypoint and line extremities detection utilized a modified HRNetV2-w48 [34] as the encoder’s backbone network. HRNetv2 [34] is a new family of convolutional networks that maintains high-resolution representations through the whole process resulting in semantically richer and spatially more precise representations. To improve the spatial resolution of the predicted heatmaps, we incorporated 2×2$\times$2 × upsampling and concatenated skip-connection features from the corresponding resolution of the convolution stem to fuse the features at different scales. Additionally, a more lightweight version of the backbone is achieved by reducing dense representation layer size in the network’s last stages. The final predictions exhibit half the resolution of the original image, with softmax and sigmoid employed as the final activation functions for keypoints and lines, respectively.
3.2.2 Keypoints Mask
@@
When homography was unavailable due to a limited number of points, the heatmaps associated with points belonging to $\mathcal{K}p_{1}$, $\mathcal{K}p_{2}$, and $\mathcal{K}p_{3}$—which would have been indicated by homography—were masked out from the loss function as long as the line to which they belong is included in the GT annotation. Additionally, when the external point required to compute ellipse tangent points in $\mathcal{K}p_{2}$ is not present in $\mathcal{K}p$, the pair of tangent points candidates is also excluded.
3.3 Camera Projection Model
We employ a standard full perspective camera model:
$\mathbf{P}=\mathbf{KR}[\mathbf{I}\,|\,\mathbf{-c}]\in\mathbb{R}^{3\times 4},$ (1)
@@ 3.3 Camera Projection Model
where $\mathbf{K}\in\mathbb{R}^{3\times 3}$ denotes the intrinsics to transform from camera coordinates to image ones, and $\mathbf{R}\in\mathbb{R}^{3\times 3}$ and $\mathbf{c}\in\mathbb{R}^{3}$ determine the extrinsics (rotation and translation) to map from scene coordinates to camera ones. Following [20], we assume zero skew and a known pixel aspect ratio. Additionally, for simplicity, we assume the principal point coincides with the center of the image, and we neglect astigmatism or distortions.
3.3.1 Camera Parameters Estimation
Extrinsic and intrinsic parameters are inferred by leveraging the coordinates of 3D object points and their corresponding 2D projections using the soccer field model as a calibration rig, following [39], which consists of a closed-form solution followed by a non-linear refinement based on the maximum likelihood criterion. We follow  [15] to calibrate the 3D soccer field model rig: we consider two additional vertical planes containing the goal polygons, including non-planar points such as keypoints belonging to the goal posts and crossbars. Additionally, we extend this strategy to enhance completeness by providing estimations when insufficient points are on the ground plane, calibrating over the vertical planes, and subsequently transforming camera parameters to the ground-plane coordinate system.
@@ 3.3.1 Camera Parameters Estimation
To account for keypoint misdetections and other complexities in camera parameter retrieval, such as frames with only one non-planar keypoint visible, the calibration process was repeated on several subsets of keypoints. Similar to  [15], these subsets were selected based on various heuristics: full-keypoints, including all keypoints sets $\mathcal{K}p$, $\mathcal{K}p_{e}$, $\mathcal{K}p_{1}$, $\mathcal{K}p_{2}$ and $\mathcal{K}p_{3}$; main-keypoints, comprising only line-line intersections from the original SoccerNet annotations [8]; and ground-plane-keypoints, which excludes non-planar keypoints. Furthermore, we applied a grid of RANSAC [16] reprojection error thresholds to each subset. The final camera calibration values were determined through a heuristic voting process, prioritizing the method yielding a lower reprojection error, with emphasis on the full-keypoints subset.
3.3.2 Homography Estimation
@@
Assuming the world coordinate system such that $z=0$ corresponds to the ground plane, the ground-to-image homography $\mathbf{H}$ can be obtained from the first, second, and fourth columns of the camera projection matrix $\mathbf{P}$. Nevertheless, inaccurate estimations for keypoints associated with the non-planar rig, such as those belonging to the goal posts and crossbars, may result in a flawed homography estimation. To address this issue, we employ classical homography estimation with DLT [20] and RANSAC on the ground-plane-keypoints subset. We define a maximum allowable reprojection error to consider a point pair as an inlier and subsequently refine the computed homography matrix using the Levenberg-Marquardt method [20] on the point correspondences and initial homography estimation.
4 Experiments

### v3

@@ III-A2 Keypoint Disambiguation
Due to the multi-view nature of the SoccerNet dataset [12] and, for instance, considering one of the soccer field’s semi-circles, ambiguity appears in its respective $\mathcal{K}p_{1}$ and $\mathcal{K}p_{2}$ keypoints candidates, as shown in Fig. 2-Bottom. This is, while the expected locations of keypoint pairs in the image are known, the challenge lies in uniquely identifying and matching each individual keypoint. To solve that, we define two different strategies to handle that disambiguation depending on the total number of keypoints generated in the previous sets: when there are sufficient points in the $\mathcal{K}p\cup\mathcal{K}e$ set to infer a homography ($\mathbf{H}$), i.e., four points, $\mathcal{K}p_{1}$ is computed first by choosing the candidates combination that minimizes the reprojection error. Then, we include $\mathcal{K}p_{1}$ to newly infer a homography estimation ($\mathbf{H_{1}}$) and repeat the same strategy on the $\mathcal{K}p_{2}$ set (obtaining $\mathbf{H_{2}}$), as it is shown in Fig. 1. Otherwise, we perform a grid-search involving both $\mathcal{K}p_{1}$ and $\mathcal{K}p_{2}$ candidates when $\mathcal{K}p\cup\mathcal{K}p_{e}\cup\mathcal{K}p_{1}^{*}\cup\mathcal{K}p_{2}^{*}\geq 4$, where * denotes a possible candidate combination. The grid-search iterates over all keypoints candidates in a set-wise manner to avoid unfeasible combinations and keeps the one with minimum reprojection error. In this case, $\mathbf{H_{1}}$ estimation is bypassed, and $\mathbf{H_{2}}$ is computed only after resolving both $\mathcal{K}p_{1}$ and $\mathcal{K}p_{2}$. It is worth pointing out that no more keypoint sets beyond $\mathcal{K}p\cup\mathcal{K}p_{e}$ are computed if neither of the above stated conditions are met. Moreover, once we compute all the pre-defined keypoint sets, two additional geometrical constraints are applied in case that homography estimation or ellipse fitting is not accurate enough. Initially, we manually establish a reprojection error threshold to validate keypoints. Subsequently, through iteration over combinations of keypoints, we construct vectors and ensure that the cross-products maintain consistent signs in both world and image coordinates. This final step is essential in cases where two distinct combinations yield valid top- and bottom-view perspectives of the field while exhibiting identical reprojection errors. Utilizing cross-products enables us to differentiate and retain the keypoint combination corresponding to the field’s top-view. The full keypoint generation process is depicted in Fig. 1-top.
@@
III-A3 Left-Right Disambiguation
In sequences where the camera angle aligns with the longitudinal axis of the court, an ambiguity arises regarding the distinction between the right and left halves of the field. Hence, a critical step to ensure consistency and robustness across keypoints and lines detection processes involves differentiating between the two sides. Taking into account the camera calibration evaluation protocol in [8], which considers both of the ambiguous options and keeps the highest score, this is accomplished by implementing a remap to the ground-truth (GT) values, ensuring that the goal area closest to the camera consistently represents the left side. The process of checking whether or not the mapping should be applied is defined in a heuristic fashion. We compute angles of horizontal and vertical soccer field lines, respectively; and then set a threshold taking into account angle distribution and visual inspection. This approach facilitates an effective model training process by deferring the disambiguation task until after the inference stage as an extra step if needed.
III-B Keypoints and Lines Detection
@@
Our approach is built upon Falaleev et al. [14] solution, which makes use of two encoder-decoder convolutional neural networks to estimate the position of the pre-defined keypoints and the soccer field lines depicted in [8] excluding conics. While in [14] the line model is given an auxiliary role to enhance keypoint completeness, in our approach is used as the key component of our refinement module in the last calibration stage. During inference, the former produces heatmaps for each pre-defined keypoint with a single $2$-pixel Gaussian peak sigma positioned in the keypoint location, accompanied by an additional background channel. This additional channel reflects the inverse of the maximum value among the other target feature maps, ensuring that the resultant target tensor behaves as a probability distribution function at every spatial point. Meanwhile, the latter network produces heatmaps for each visible soccer field line within the frame, assigning two Gaussian peaks at the line extremities’ locations. Additionally, we introduce an extra channel, known as the boundary channel, to our heatmap following the approach outlined in [50]. This augmentation aims to enhance the efficient capture of global information regarding the soccer field and improve extremities detection, particularly near image borders. We effectively extract the position of keypoints and line extremities from the generated heatmaps by employing a max pooling operation, which calculates the maximum value for patches of a feature map, drawing inspiration from the methodology proposed in [55]. This process is summarized on Fig. 1-Bottom.
III-B1 Architecture
@@
Following [14], the keypoint and line extremities detection utilized a modified HRNetV2-w48 [49] network as the encoder’s backbone network. HRNetv2 [49] is a new family of convolutional networks that maintains high-resolution representations through the whole process resulting in semantically richer and spatially more precise representations. To improve the spatial resolution of the predicted heatmaps, we incorporated 2×2$\times$2 × upsampling and concatenated skip-connection features from the corresponding resolution of the convolution stem to fuse the features at different scales. The resulting feature maps are produced at half the spatial resolution of the input image, striking a balance between computational efficiency and spatial detail retention. Additionally, a more lightweight version of the backbone is achieved by reducing the dense representation layers sizes in the network’s last stages. The final predictions exhibit half the resolution of the original image, with softmax and sigmoid employed as the final activation functions for keypoints and lines, respectively.
III-B2 Keypoints Mask
@@
When homography was unavailable due to a limited number of points, the heatmaps associated with points belonging to $\mathcal{K}p_{1}$, $\mathcal{K}p_{2}$, and $\mathcal{K}p_{3}$—which would have been derived from the homography itself—were masked out from the loss function as long as the line to which they belong is included in the GT annotation. Additionally, when the external point required to compute ellipse tangent points in $\mathcal{K}p_{2}$ is not present in $\mathcal{K}p$, the pair of tangent points candidates is also masked out.
III-C Camera Projection Model
We employ a standard full-perspective camera model as:
$\mathbf{P}=\mathbf{KR}[\mathbf{I}\,|\,\mathbf{-t}]\in\mathbb{R}^{3\times 4},$ (1)
where $\mathbf{R}\in\mathbb{R}^{3\times 3}$ and $\mathbf{t}\in\mathbb{R}^{3}$ denote the extrinsic parameters (rotation and translation, respectively) to map from scene coordinates to camera ones; and $\mathbf{K}\in\mathbb{R}^{3\times 3}$, which includes the intrinsic parameters to transform from camera coordinates to image ones. The latter has the form:
$\mathbf{K}=\begin{bmatrix}\alpha_{x}&s&x_{0}\\ &\alpha_{y}&y_{0}\\ &&1\end{bmatrix},$ (2)
@@ III-C Camera Projection Model
where $\{\alpha_{x},\alpha_{y}\}$ denote the focal length, $\{x_{0},y_{0}\}$ the principal point, and $s$ the skew value. Following [22], we assume zero skew and a known pixel aspect ratio. Additionally, for simplicity, we assume the principal point ($x_{0},y_{0}$) coincides with the center of the image, i.e., we neglect astigmatism or distortions.
III-C1 Camera Parameters Estimation
@@
Extrinsic and intrinsic parameters in Eq. (1) are inferred by leveraging the coordinates of 3D object points and their corresponding 2D projections using the soccer field model as a calibration rig, following [54], which consists of a closed-form solution followed by a non-linear refinement based on the maximum likelihood criterion. To calibrate the 3D soccer field model rig, we consider two additional vertical planes containing the goal polygons, including non-planar points such as keypoints belonging to the goal posts and crossbars. Additionally, we extend this strategy to enhance completeness by providing estimations when insufficient points are on the ground plane, calibrating over the vertical planes, and subsequently transforming camera parameters to the ground-plane coordinate system. When a sufficient number of points, i.e., 6 keypoints, is available, an initial estimate of the camera’s intrinsic matrix $\mathbf{K}$ is inferred. Camera calibration is then performed using this estimation, resulting in a more robust and stable calibration. Otherwise, calibration is done by jointly optimizing all three unknowns: $\mathbf{K}$, $\mathbf{R}$, and $\mathbf{t}$. To account for keypoint misdetections and other complexities in camera parameter retrieval, such as frames with only one non-planar keypoint visible, the calibration process was repeated on several subsets of keypoints. Similar to [14], these subsets were selected based on various heuristics: full-keypoints, including all keypoints sets $\mathcal{K}p$, $\mathcal{K}p_{e}$, $\mathcal{K}p_{1}$, $\mathcal{K}p_{2}$ and $\mathcal{K}p_{3}$; main-keypoints, comprising only line-line intersections from the original SoccerNet annotations [8]; and ground-plane-keypoints, which excludes non-planar keypoints. Furthermore, we applied a grid of RANSAC [15] reprojection error thresholds to each subset. The final camera calibration values were determined through a heuristic voting process, prioritizing the method yielding a lower reprojection error, with emphasis on the full-keypoints subset.
III-C2 Homography Estimation
Assuming the world coordinate system such that $z=0$ corresponds to the ground plane, the ground-to-image homography $\mathbf{H}$ can be obtained from the first, second, and fourth columns of the camera projection matrix $\mathbf{P}$ as:
$\mathbf{H}\cong\mathbf{KR}\begin{bmatrix}1&0&-t_{x}\\ 0&1&-t_{y}\\ 0&0&-t_{z}\end{bmatrix}\in\mathbb{R}^{3\times 3},$ (3)
@@ III-C2 Homography Estimation
where $\mathbf{t}=[t_{x},t_{y},t_{z}]^{\top}$. Nevertheless, inaccurate estimations for keypoints associated with the non-planar rig, such as those belonging to the goalposts and crossbars, may result in a flawed homography estimation. To address this issue, we employ classical homography estimation with DLT [22] and RANSAC [15] on the ground-plane-keypoints subset. We define a maximum allowable reprojection error to consider a point pair as an inlier and subsequently refine the initial homography estimation matrix using the Levenberg-Marquardt method [22] on the 2D-3D point correspondences. Similarly to the approach used for camera parameter estimation, we applied a grid search over RANSAC reprojection error and employed the heuristic voting method, but in this instance, it is restricted to the ground-plane-keypoints subset.
Figure 3: PnL refinement module. Top-Left: Let $\{\mathbf{p_{d}},\mathbf{q_{d}}\}$ be the visible extremities of a detected 3D soccer field line and $\{\mathbf{\bar{p}_{d}},\mathbf{\bar{q}_{d}}\}$ their projected 2D points in the image plane, which define the projected line $\mathbf{l_{d}}$ as well as the detected line projection given an initial calibration estimate. Top-Right: $d_{1}$ and $d_{2}$ represent the line-based reprojection error between $\mathbf{l_{d}}$ (in red) and $\mathbf{l}$ (in blue). Bottom: Soccer field’s top-view. Estimated camera pose $\{\mathbf{R},\mathbf{t}\}$ defines the camera plane $\mathbf{\Pi_{c}}$. Let $\{\mathbf{p},\mathbf{q}\}$ be the real extremities of a known 3D line given a field model, and $\mathbf{q_{c}}$ the corrected 3D point through the line-camera plane intersection. Note that $\mathbf{p}\equiv\mathbf{p_{d}}$ since the real line extremity is visible in the camera frame. The green area corresponds to the camera viewing cone.
III-D Point and Line Refinement Module
To address the sparsity of keypoints on the grid, particularly in regions distant from the penalty boxes, where the majority of keypoints are concentrated, and to mitigate issues arising from keypoint misdetections overlapping with detected lines, we have developed a refinement module which makes use of the detected lines information. This module, termed the Point and Line (PnL) refinement module, enhances both homography and calibration estimates by jointly leveraging the information from detected keypoints and lines. The PnL module integrates line information to complement the keypoint data, providing a more robust and comprehensive basis for accurate field registration and camera calibration, especially in areas where keypoint information alone may be insufficient or unreliable. Setting the internal camera calibration matrix $\mathbf{K}$ as a fixed term, PnL refinement module will optimize the calibration estimate through the camera pose $\bm{\Theta}=\{\mathbf{R,t}\}$ space, which includes the rotation and translation parameters, respectively.
Inspired by previous approaches [36, 48, 17], we next describe the line parameterization, the error function, and its integration within the PnL refinement module. Let $\mathbf{p},\mathbf{q}\in\mathbb{R}^{3}$ represent the real extremities of a known 3D soccer field line based on a field model, and let $\mathbf{p_{d}},\mathbf{q_{d}}\in\mathbb{R}^{3}$ denote the visible extremities of the same 3D soccer field line detected in the image frame, with image coordinates $\bar{\mathbf{p}}_{d},\bar{\mathbf{q}}_{d}\in\mathbb{R}^{2}$ defining the detected soccer field line $\mathbf{l}_{d}=\overrightarrow{\bar{\mathbf{p}}_{d}\bar{\mathbf{q}}_{d}}$. Note that if the real line extremities are visible in the camera frame, they are equivalent to the detected extremities, i.e., $\mathbf{p}\equiv\mathbf{p_{d}}$. Otherwise, the 3D positions of $\mathbf{p_{d}}$ and $\mathbf{q_{d}}$ are unknown.
By projecting $\mathbf{p}$ and $\mathbf{q}$ into the image plane, given an initial calibration estimate $\bm{\Theta}$ (or a homography matrix $\mathbf{H}$ in the case of planar estimation), we define the projected soccer field line $\mathbf{l}$ and compute both $\bar{\mathbf{p}}$ and $\bar{\mathbf{q}}$ as its intersections with the image boundaries (see Fig. 3-Top). Hence, analogously to the detected lines, $\mathbf{l}=\overrightarrow{\bar{\mathbf{p}}\bar{\mathbf{q}}}$.
However, it is worth noting that 3D line extremities $\mathbf{p},\mathbf{q}$ lying behind the field’s camera plane $\mathbf{\Pi_{c}}$—i.e., points whose image projections have $z_{c}<0$—cannot be projected. To address this issue, we construct the field’s camera plane using the estimated camera position $\mathbf{t}$ and a normal vector $\mathbf{n_{t}}$ to the camera plane (see Fig. 3-Bottom). The latter is built based on the camera position and the projection of the principal point onto the planar field. In the case of planar estimation, camera position and rotation are obtained directly from homography matrix [22]. Therefore, to correctly project $\mathbf{l}$ onto the image, $\mathbf{p}$ and $\mathbf{q}$ must be checked and corrected to satisfy the next planar condition:
$\mathbf{q_{c}}=\begin{dcases}\mathbf{q}+\frac{\mathbf{n_{t}}\cdot(\mathbf{t}-\mathbf{q})}{\mathbf{n_{t}}\cdot(\mathbf{p}-\mathbf{q})}+\epsilon(\mathbf{n_{t}})&\text{if $(\mathbf{q}-\mathbf{t})\cdot\mathbf{n_{t}}<0$}\\ \mathbf{q}&\text{otherwise}\end{dcases},$ (4)
with $\epsilon(\mathbf{n_{t}})$ being a threshold added in the direction of the normal vector to ensure that $\mathbf{q_{c}}$ lies ahead of the camera plane $\mathbf{\Pi_{c}}$. $\cdot$” denotes a dot product. Once the projected extremities are corrected, the line $\mathbf{l}$, and consequently $\bar{\mathbf{p}}$ and $\bar{\mathbf{q}}$, can be computed. Given the image observations $\bar{\mathbf{p}}_{d}$ and $\bar{\mathbf{q}}_{d}$, we then define the point-to-line distances, following [43], to the projected line $\mathbf{l}$ as follows:
[…]
where $\Delta_{y}=\bar{\mathbf{q}}_{y}-\bar{\mathbf{p}}_{y}$ and $\Delta_{x}=\bar{\mathbf{q}}_{x}-\bar{\mathbf{p}}_{x}$. It is important to note that by representing lines using their endpoints, we obtain comparable error representations for both points and lines. Given an image detection $\bar{\mathbf{x}}_{d}$ of a predefined keypoint, we define a reprojection error as:
$d(\bar{\mathbf{x}}_{d},\mathbf{x})=\bar{\mathbf{x}}_{d}-\bm{\pi}(\bm{\Theta},\mathbf{x}),$ (6)
where $\bm{\pi}(\bm{\Theta},\mathbf{x})$ represents the projection of the keypoint world coordinate $\mathbf{x}\in\mathbb{R}^{3}$ onto the image plane, given a calibration estimate $\bm{\Theta}$. We can thus construct a unified cost function that integrates each of the error terms as follows:
$C=\alpha\sum_{i\in\mathcal{L}}d(\mathbf{l}_{i},\mathbf{l}_{d,i})+(1-\alpha)\sum_{j\in\mathcal{K}p_{c}}d(\bar{\mathbf{x}}_{d,j},\mathbf{x}_{j}),$ (7)
where $\mathcal{L}$ represents the set of detected lines and $\mathcal{K}p_{c}=\mathcal{K}p\cup\mathcal{K}p_{e}\cup\mathcal{K}p_{1}\cup\mathcal{K}p_{2}\cup\mathcal{K}p_{3}$ denotes the set-theoretic union of all available keypoint sets. A weighting parameter, $\alpha$, is empirically set to balance the influence of points and lines in the cost function. Following [36], a recursive approach over the detected reprojection line and point error will be applied to optimize the pose parameters $\bm{\Theta}$ as a nonlinear least-squares problem. To mitigate line misdetection issues, each detected line is compared to its projected counterpart using the initial estimate $\bm{\Theta}$. A projection error threshold is applied to both extremities of the detected line before it is subsequently added to the $\mathcal{L}$ set.
IV Experiments